# Cryoscope: A Cryogenic Infrared Survey Telescope

### v1

@@ 5 Summary and Path Forward
To this end, we are developing a new telescope that will provide the coverage needed for discoveries in dynamic infrared astronomy. Cryoscope uses a novel optomechanical design to keep its optical path at cryogenic temperatures, eliminating thermal background from the ground within the K-dark bandpass (2.25-2.5 $\,\mu$m). This allows the instantaneous field-of-view to be increased 80×80$\times$80 × that of other previous deep surveyors in the K-band (e.g., VISTA), at no cost to its sensitivity. To image its large focal plane, Cryoscope will employ a new infrared camera with more than 600 Mpix, the highest number of pixels ever made for infrared astronomy. To achieve the depth demanded for the greatest variety of science cases, Cryoscope will be uniquely situated at Dome C, Antarctica, relying upon infrared skies that are at least 30 times darker than the best astronomical sites at temperate latitudes (e.g., Mauna Kea, Cerro Pachón). Dome C is already home to the productive ASTEP+ project, which has complemented TESS to propel exoplanet science in the optical bands. We now seek Dome C’s infrared sky. Reduced sky brightness in K-dark over the elevated Antarctic plateau, coupled with clear weather, unparalleled seeing, and continuously dark winter nights, makes the location of Dome C critical for any ground-based, high-cadence, deep and wide infrared surveyor.
@@
We have built a fifth-scale prototype of the full-scale telescope, the Cryoscope Pathfinder, to demonstrate that these obstacles are surmountable and these science cases are within reach. Due to be deployed in December 2025, the Pathfinder will for the first time measure the K-dark sky brightness at Dome C during the winter observing season and retire the technological risk of operating a fully cryogenic telescope in the Antarctic. Cryoscope Pathfinder is currently being tested in the lab and has been shown to deliver diffraction-limited imaging across its field-of-view. Upcoming tests will soon validate the thermal management strategy to minimize the telescope thermal emission well below the dark sky levels conservatively predicted at Dome C. On-sky tests will then demonstrate focus stability and verify the image quality revealed in the lab.
@@ Appendix A Volumetric survey speed across the UVOIR
Cryoscope Pathfinder Planned 2025 16.2 8$\times$(10+5) $K_{dark}$ 18.0 $1.80\times 10^{2}$
@@
References. — UVEX: Table 1, Kulkarni et al. (2021). ULTRASAT: Table 1, Shvartzvald et al. (2024). GALEX AIS: Table 1, Martin et al. (2005). Rubin: Table 2, Bianco et al. (2022), using science-required limited depths. ZTF: Bellm et al. (2019). Roman: Two reference surveys are considered, the high latitude time domain survey (HLTDS) and Galactic bulge time domain survey (GBTDS). A selection of the reddest filters are chosen from HLTDS in Rose et al. (2021). $10\sigma$ depths presented in Table 2 are adjusted to $5\sigma$ depths for the same exposure time, assuming background-limited observations: $m_{5\sigma}\approx m_{10\sigma}+2.5\log 2$. For proposed concepts of GBTDS, the nominal survey cycles through $\sim$ 2 deg2 total area every 15 min. We assume $15/(2/0.281)\sim 2$ min visits including overheads with nominal 73 s exposures in two proposed filter combinations. We approximate the limiting depth reached in these stares in comparison to depths accomplished in 57 s background-limited observations: $m_{73\rm{s}}\approx m_{57\rm{s}}+1.25\log(73/57)$. See Roman Core Community Survey (2025) for preliminary details for GBTDS and Roman WFI and Observatory Performance (2024) for depths in 57 s single exposures. Euclid: See Euclid Collaboration et al. (2022), Section 4 for the breakdown between overhead and exposure time for a given visit in the reference observing survey (ROS) and Table 7 for $5\sigma$ depths. Only considering NISP and the depths reached therein, yet across the entire NISP FoV instead of the shared NISP+VIS FoV (0.55 instead of 0.53 deg2). Palomar Gattini-IR: De et al. (2020). WINTER: Internal communication at Caltech, nightly survey operations currently in J-band. VISTA: McMahon et al. (2013); Sutherland et al. (2015), Active field of view is 0.6 deg2. Cryoscope+Pathfinder: This work. WISE: Sensitivities in Wright et al. (2010) correspond to co-added stacks of 8 exposures. We define the 8 exposures to encompass a visit.
@@ Appendix B First principles argument for a cryogenic optical path
However, assuming blackbody emission from the sky is a crude approximation. In the sensitivity estimates in the main text, we find the power delivered to the detector from the thermal sky by integrating over the South Pole sky spectrum measured by Ashley et al. (1996) and Nguyen et al. (1996), a conservative upper estimate for Dome C. Integrating over the sky spectrum, we directly compute the signal from the sky and compare to the estimated thermal emission from the telescope walls. The thermal background at the central detector pixel is shown in the left panel of Figure 20 for a range of cryostat temperatures. The contribution from the South Pole thermal sky alongside fiducial constant sky brightness curves are overplotted. Only below $\sim-80^{\circ}$ C does thermal emission from the cryostat walls become sub-dominant to the South Pole sky brightness within the K-dark bandpass. We show in the right panel of Figure 20, how the Pathfinder’s maximum attainable survey speed depends on its interior wall temperature under skies with brightnesses of 10, 100, and 1000 $\mu$Jy/arcsec2. As expected, bright skies fundamentally limit sensitivity, and a cryogenic optical path provides minimal gains towards increasing survey speed. However, as skies become darker, a cryogenic optical path becomes relevant. Limiting depths under the darkest skies can be increased by two or more magnitudes and survey speed enhanced five to tenfold compared to those under nominal ambient conditions by maintaining the optics core at cryogenic temperatures. In the K-dark window at Dome C, thermal self-emission can no longer be considered sub-dominant to the thermal sky. Cryocooling ensures that regardless of variations in Dome C sky brightness or ambient temperature fluctuations at ground level, survey speed is always maximized. The sky background, instead of the telescope’s own thermal emission, is the fundamental dominant thermal background.
Figure 20: Left: Estimated incident thermal background at the central detector pixel. Blackbody thermal self-emission from the cryostat walls is estimated as prescribed in the text from $-20^{\circ}$ to $-80^{\circ}$ C, marked by colored dashed lines. Fiducial constant sky brightness curves are over-plotted in gray, alongside the measured South Pole sky background from Ashley et al. (1996) and Nguyen et al. (1996) and Cryoscope’s K-dark filter response curve. Right: Maximum attainable volumetric survey speed for the Pathfinder as a function of the temperature of the cryostat walls, assuming the fiducial values for K-dark sky brightness in the left panel.
Appendix C Conversion between AB and Vega magnitudes in K-dark
We derive the isophotal flux density and AB magnitude for Vega in the K-dark filter for the Pathfinder and full-scale telescope. From Eq. 9 in Tokunaga & Vacca (2005),
@@ Appendix C Conversion between AB and Vega magnitudes in K-dark
The isophotal flux density is defined in Eq. 6 by Tokunaga & Vacca (2005), integrating over the frequency interval of the K-dark filter.
@@
where $\nu_{\rm iso}$ is the isophotal frequency and $S(\nu)$ is the total system response: $S(\nu)=T(\nu)Q(\nu)R(\nu)A_{\rm tel}$, with $T$ being the atmospheric transmission, $Q$ the product of the total optical throughput and quantum efficiency, $R$ the filter response, and $A_{\rm tel}$ the telescope collecting area. As an estimate for the conversion from K-dark AB magnitudes to Vega magnitudes, we presume the sky transmission profile shown in Figure 3 in the main text. The optical throughput of the telescope, incorporating all losses, is 0.83. The QEs for the H2RG and SATIN detector were incorporated for the Pathfinder and full-scale, respectively, and the collecting areas for each telescope fall out of Eq. C2. We use the Vega spectrum in $\tt{alpha\_lyr\_stis\_010.fits}$ distributed by the synphot package (STScI Development Team, 2018).
The conversion between AB and Vega in K-dark is estimated to be:
@@
Table 5: K-dark isophotal wavelengths, frequencies, flux densities and AB magnitudes for Vega

### v2

@@ 4.8 High-energy neutrino counterparts
4.9 Supermassive Black Holes
Sitting at the heart of nearly every galaxy is a supermassive black hole (SMBH) that can be millions to billions times the mass of the Sun. These reveal themselves across the electromagnetic spectrum through interactions with stars and gas in their environment: when a star passes within the sphere of disruption of the SMBH, tidal forces will tear the star apart with a flash of radiation known as a tidal distruption event (TDE; Rees, 1988), while gas funneling in can form an optically thick, geometrically thin accretion disk around the SMBH creating an active galactic nuclei (AGN). SMBHs can be energetically dominant in their host galaxies and even dictate how the galaxy evolves. Variability is a key characteristic of SMBHs and has been extensively studied in optical and X-ray wavelengths, but IR studies have been largely limited to the six-month cadence decadal data set that WISE provided at mid-infrared wavelengths. IR variability probes the outer regions of the accretion disk and the dusty torus around the SMBH and can feature significant contribution from the host galaxies, providing important insights into these environments.
While the vast majority of TDEs to date have been discovered based on their X-ray (Saxton et al., 2020) or UV/optical emission (van Velzen et al., 2020), IR wavelengths offer critical diagnostics that cannot be gleaned from the current sample. In particular, Cryoscope observations will be exquisitely sensitive to the presence of dust echoes from the initial X-ray/UV emission (van Velzen et al., 2021)—in fact the Cryoscope $K_{dark}$ bandpass directly samples the peak of thermal emission expected for the dust heating ($T\sim 1800$  K). Based only on the observed sample of UV/optical TDE discoveries, with a 10,000  deg2 survey starting when the direct drive becomes available in year 2, we expect Cryoscope Pathfinder to discover $\sim$ a dozen new IR-selected TDEs each year. By comparing the properties of the IR dust emission with the UV/optical emission, we can directly measure both the dust location (based on the lag of the IR signal) as well as the dust-covering fraction (based on the ratio of IR to UV/optical luminosity; Jiang et al., 2021). Such measurements are exceedingly difficult to obtain by any other means, providing a completely unique picture of the geometry of the SMBH environment in distant, quiescent galaxies.
Several hundred AGN have been identified in the past decade that exhibit both strong photometric and spectroscopic optical variability known as changing-state AGN (CS-AGN; Ricci & Trakhtenbrot, 2023). The physical mechanism is not clear but seems to be related to an accelerating outflow or varying accretion rates onto the SMBH, which may in turn be driven by physical changes in the accretion disk or its magnetic field support. IR variability is seen to echo the optical variability with large amplitude variations and a time lag expected for dust reprocessing. Cryoscope enables novel studies of CS-AGN with much higher cadences. Cryoscope will be able to probe the interface region between the accretion disk and the torus; identify higher redshift CS-AGN where rest-frame optical continuum changes have been redshifted into the observed near-IR frame; and detect near-IR CS-AGN where the optical variability may be obscured by dust. There is also the possibility of capturing CS-AGN where a triggering event occurs on the outskirts of the accretion disk and a thermal front then propagates radially inwards, temporarily altering the disk geometry.
Supermassive black hole binaries (SMBHBs) are an expected consequence of hierarchical models of galaxy formation and an important source of nanohertz frequency gravitational waves (detectable by LISA). At close separations, they are not physically resolvable, but induced effects from the kinematics of the rotating pair may be imprinted on electromagnetic (EM) radiation from the SMBHB. The two main physical mechanisms for an EM signature are: (i) relativistic Doppler boosting of emission from a mini-disk associated with one of the binary; and (ii) periodic modulation of the accretion rate on the SMBHB. A significant distinction between the two is that variability due to accretion rate fluctuations should be isotropic, whereas relativistic Doppler modulation produces anisotropic variability resembling a rotating forward-beamed lighthouse surrounded by fog. This can have observational consequences for IR variability depending on the geometry of the AGN system, e.g., the ratio of dust light crossing time to the source variability period, torus inclination, and opening angle. In particular, there should be a population of nearly face-on binaries with an inclined dust torus and modulated IR variability but not observable optical/UV variability that would have been missed in optical searches, but which Cryoscope has the sensitivity and field coverage to easily find.
Finally, the most powerful technique to probe AGN is reverberation mapping (RM), where measured light travel time delays map out spatial scales that are often well beyond imaging capabilities of current telescopes. The RM technique was pioneered in the optical where the location and motion of gas flows that constitute the broad-line region are constrained by observing the time-delayed response of broad emission lines to the variable continuum flux from the accretion disk. Near-IR RM tracks emission produced in the dusty torus: the dust is irradiated by X-ray and UV/optical continuum emission from the inner regions and is reprocessed as thermal emission in the $1300-2000$ K range. The dust only survives at temperatures below the sublimation temperature, and therefore this temperature defines the inner radius of the dusty torus. As the torus is much larger and further out than the inner accretion disk, the IR variability is longer and weaker, and the corresponding reverberation lags are longer. Cryoscope again provides the capability for extensive near-IR monitoring of the AGN population, which coupled with optical light curves from Rubin will create an unprecedented sample of optical-to-IR lags over a range of black hole masses, accretion rates, and orientations.
@@ 5 Summary and Path Forward
To this end, we are developing a new telescope that will provide the coverage needed for discoveries in dynamic infrared astronomy. Cryoscope uses a novel optomechanical design to keep its optical path at cryogenic temperatures, eliminating thermal background from the ground within the $K_{dark}$ bandpass (2.25-2.5 $\,\mu$m). This allows the instantaneous field-of-view to be increased 80×80$\times$80 × that of other previous deep surveyors in the $K$-band (e.g., VISTA), at no cost to its sensitivity. To image its large focal plane, Cryoscope will employ a new infrared camera with more than 600 Mpix, the highest number of pixels ever made for infrared astronomy. To achieve the depth demanded for the greatest variety of science cases, Cryoscope will be uniquely situated at Dome C, Antarctica, relying upon infrared skies that are at least 30 times darker than the best astronomical sites at temperate latitudes (e.g., Mauna Kea, Cerro Pachón). Dome C is already home to the productive ASTEP+ project, which has complemented TESS to propel exoplanet science in the optical bands. We now seek Dome C’s infrared sky. Reduced sky brightness in $K_{dark}$ over the elevated Antarctic plateau, coupled with clear weather, unparalleled seeing, and continuously dark winter nights, makes the location of Dome C critical for any ground-based, high-cadence, deep and wide infrared surveyor.
@@
We have built a fifth-scale prototype of the full-scale telescope, the Cryoscope Pathfinder, to demonstrate that these obstacles are surmountable and these science cases are within reach. Due to be deployed in December 2026, the Pathfinder will for the first time measure the $K_{dark}$ sky brightness at Dome C during the winter observing season and retire the technological risk of operating a fully cryogenic telescope in the Antarctic. Cryoscope Pathfinder is currently being tested in the lab and has been shown to deliver diffraction-limited imaging across its field-of-view. Upcoming tests will soon validate the thermal management strategy to minimize the telescope thermal emission well below the dark sky levels conservatively predicted at Dome C. On-sky tests will then demonstrate focus stability and verify the image quality revealed in the lab.
@@ Appendix A Volumetric survey speed across the UVOIR
SPHEREx All-sky Survey First light 2025 12.25 112 Band 1 (0.75–1.11 $\,\mu$m) 19.517 $9.15\times 10^{2}$
Band 2 (1.11–1.64 $\,\mu$m) 19.522 $9.20\times 10^{2}$
Band 3 (1.64–2.42 $\,\mu$m) 19.6 $9.73\times 10^{2}$
Band 4 (2.42–3.82 $\,\mu$m) 19.8 $1.23\times 10^{3}$
Band 5 (3.82–4.42 $\,\mu$m) 18.7 $3.30\times 10^{2}$
Band 6 (4.42–5 $\,\mu$m) 18.1 $1.50\times 10^{2}$
Cryoscope Pathfinder Planned 2026 16.2 8$\times$(10+5) $K_{dark}$ 18.0 $1.80\times 10^{2}$
@@
NEO Surveyor Planned 2027 11.89 6$\times$30 NC1 (4–5.2 $\,\mu$m) 19.4 $4.61\times 10^{2}$
NC2 (6–10 $\,\mu$m) 18.8 $2.29\times 10^{2}$
References. — UVEX: Table 1, Kulkarni et al. (2021). ULTRASAT: Table 1, Shvartzvald et al. (2024). GALEX AIS: Table 1, Martin et al. (2005). Rubin: Table 2, Bianco et al. (2022), using science-required limited depths. ZTF: Bellm et al. (2019). Roman: Two reference surveys are considered, the high latitude time domain survey (HLTDS) and Galactic bulge time domain survey (GBTDS). A selection of the reddest filters are chosen from HLTDS in Rose et al. (2021). $10\sigma$ depths presented in Table 2 are adjusted to $5\sigma$ depths for the same exposure time, assuming background-limited observations: $m_{5\sigma}\approx m_{10\sigma}+2.5\log 2$. For proposed concepts of GBTDS, the nominal survey cycles through $\sim$ 2  deg2 total area every 15 min. We assume $15/(2/0.281)\sim 2$ min visits including overheads with nominal 73 s exposures. The primary filter to be used will be the broadband F164, with supplemental coverage via F087 and F213 at different cadences. Filter combinations have not been finalized (https://asd.gsfc.nasa.gov/roman/comm_forum/forum_17/Core_Community_Survey_Reports-rev03-compressed.pdf). We approximate the limiting depth reached in these stares in comparison to depths accomplished in 57 s background-limited observations: $m_{73\rm{s}}\approx m_{57\rm{s}}+1.25\log(73/57)$ (for depths in 57 s single exposures, see https://roman.gsfc.nasa.gov/science/WFI_technical.html). Euclid: See Euclid Collaboration et al. (2022), Section 4 for the breakdown between overhead and exposure time for a given visit in the reference observing survey (ROS) and Table 7 for $5\sigma$ depths. Only considering NISP and the depths reached therein, yet across the entire NISP FoV instead of the shared NISP+VIS FoV (0.55 instead of 0.53  deg2). Palomar Gattini-IR: De et al. (2020). WINTER: Internal communication at Caltech, nightly survey operations currently in $J$-band. VISTA: McMahon et al. (2013); Sutherland et al. (2015), Active field of view is 0.6  deg2. SPHEREx: Given SPHEREx’s linear-variable filters (LVF), we employ the median depth and survey speed across the wavelength coverage in each of the six detector bands. Each of the six detector bands, with $3.5^{\circ}\times 3.5^{\circ}$ field-of-view, is divided into 17 spectral channels (Hui et al., 2024). All-sky point-source sensitivities in each spectral channel are obtained from SPHEREx public products (https://github.com/SPHEREx/Public-products). Cryoscope+Pathfinder: This work. WISE: Sensitivities in Wright et al. (2010) correspond to co-added stacks of 8 exposures. We define the 8 exposures to encompass a visit. NEO Surveyor: Assuming monochromatic flux densities of 65 and 110  $\mu$Jy (Mainzer et al., 2023).
@@ Appendix B First principles argument for a cryogenic optical path
However, assuming blackbody emission from the sky is a crude approximation. In the sensitivity estimates in the main text, we find the power delivered to the detector from the thermal sky by integrating over the South Pole sky spectrum measured by Ashley et al. (1996) and Nguyen et al. (1996), a conservative upper estimate for Dome C. Integrating over the sky spectrum, we directly compute the signal from the sky and compare to the estimated thermal emission from the telescope walls. The thermal background at the central detector pixel is shown in the left panel of Figure 19 for a range of cryostat temperatures. The contribution from the South Pole thermal sky alongside fiducial constant sky brightness curves are overplotted. Only below $\sim-80$ C does thermal emission from the cryostat walls become sub-dominant to the South Pole sky brightness within the $K_{dark}$ bandpass. We show in the right panel of Figure 19, how the Pathfinder’s maximum attainable survey speed depends on its interior wall temperature under skies with brightnesses of 10, 100, and 1000  $\mu$Jy/arcsec2. As expected, bright skies fundamentally limit sensitivity, and a cryogenic optical path provides minimal gains towards increasing survey speed. However, as skies become darker, a cryogenic optical path becomes relevant. Limiting depths under the darkest skies can be increased by two or more magnitudes and survey speed enhanced five to tenfold compared to those under nominal ambient conditions by maintaining the optics core at cryogenic temperatures. In the $K_{dark}$ window at Dome C, thermal self-emission can no longer be considered sub-dominant to the thermal sky. Cryocooling ensures that regardless of variations in Dome C sky brightness or ambient temperature fluctuations at ground level, survey speed is always maximized. The sky background, instead of the telescope’s own thermal emission, is the fundamental dominant thermal background.
Figure 19: Left: Estimated incident thermal background at the central detector pixel. Blackbody thermal self-emission from the cryostat walls is estimated as prescribed in the text from $-20^{\circ}$ to $-80$ C, marked by colored dashed lines. Fiducial constant sky brightness curves are over-plotted in gray, alongside the measured South Pole sky background from Ashley et al. (1996) and Nguyen et al. (1996) and Cryoscope’s $K_{dark}$ filter response curve. Right: Maximum attainable volumetric survey speed for the Pathfinder as a function of the temperature of the cryostat walls, assuming the fiducial values for $K_{dark}$ sky brightness in the left panel.
Appendix C Conversion between AB and Vega magnitudes in $K_{dark}$
We derive the isophotal flux density and AB magnitude for Vega in the $K_{dark}$ filter for the Pathfinder and full-scale telescope. From Eq. 9 in Tokunaga & Vacca (2005),
@@ Appendix C Conversion between AB and Vega magnitudes in $K_{dark}$
The isophotal flux density is defined in Eq. 6 by Tokunaga & Vacca (2005), integrating over the frequency interval of the $K_{dark}$ filter.
@@
where $\nu_{\rm iso}$ is the isophotal frequency and $S(\nu)$ is the total system response: $S(\nu)=T(\nu)Q(\nu)R(\nu)A_{\rm tel}$, with $T$ being the atmospheric transmission, $Q$ the product of the total optical throughput and quantum efficiency, $R$ the filter response, and $A_{\rm tel}$ the telescope collecting area. As an estimate for the conversion from $K_{dark}$ AB magnitudes to Vega magnitudes, we presume the sky transmission profile shown in Figure 3 in the main text. The optical throughput of the telescope, incorporating all losses, is 0.83. The QEs for the H2RG and SATIN detector were incorporated for the Pathfinder and full-scale, respectively, and the collecting areas for each telescope fall out of Eq. C2. We use the Vega spectrum in $\tt{alpha\_lyr\_stis\_010.fits}$ distributed by the synphot package (STScI Development Team, 2018).
The conversion between AB and Vega in $K_{dark}$ is estimated to be:
@@
Table 4: $K_{dark}$ isophotal wavelengths, frequencies, flux densities and AB magnitudes for Vega